OpenStudy (anonymous):

Find all x and y (0 ≤ x, y < 2π) so that the following equation is true. (For each x and y enter answers from smallest to largest x-value.) cos (x) + i sin (y) = sin (x) + i

OpenStudy (fibonaccichick666):

so can you set the real parts and the imaginary parts equal to each other?

OpenStudy (anonymous):

ok so here. I separated real numbers from imaginary... like text tells us to. i got... cos(x)=sin(x) and sin(y)=1.... so then i thought x would be sqrt2/2 and y = 1... and its not working for me

OpenStudy (anonymous):

yes... thats what i just wrote above

OpenStudy (fibonaccichick666):

hmm, well, sine of what number =1?

OpenStudy (anonymous):

90 degrees?

OpenStudy (fibonaccichick666):

in terms of pi?

OpenStudy (anonymous):

pi/2

OpenStudy (fibonaccichick666):

cool, so is there any other way that the sin y=1 other than y=pi/2?

OpenStudy (fibonaccichick666):

on 0<y<2pi

OpenStudy (anonymous):

well no bc 270 degrees = -1

OpenStudy (fibonaccichick666):

good, so yes that is the only possible y value

OpenStudy (fibonaccichick666):

now we need sin x= cos x

OpenStudy (anonymous):

so isnt that sqrt2/2 which is pi/4

OpenStudy (fibonaccichick666):

so if you graph sine and cosine on 0<x<2pi, where do the graphs intersect?

OpenStudy (fibonaccichick666):

I am willing to bet they intersect more than once

OpenStudy (fibonaccichick666):

ok, so sqrt 2/2 is 1/sqrt 2 yes? but that isn't an angle so the sine cannot be that

OpenStudy (anonymous):

3pi/4

OpenStudy (fibonaccichick666):

try graphing, it makes life way easier

OpenStudy (freckles):

hint: sin and cos can have the same signs in quadrant 1 and also in ?

OpenStudy (fibonaccichick666):

nice hint choice

OpenStudy (anonymous):

yeah so 3pi/4 or is that wrong?

OpenStudy (fibonaccichick666):

which quadrant is 3pi/4 in?

OpenStudy (anonymous):

3

OpenStudy (fibonaccichick666):

let me see first, but here is y preference http://www.wolframalpha.com/input/?i=sinx%3Dcosx+solve+for+x&lk=4&num=1

OpenStudy (fibonaccichick666):

so 3pi/4 isn't in quadrant 3

OpenStudy (fibonaccichick666):

because it is less than pi which is less than 180

OpenStudy (fibonaccichick666):

but I agree something in quadrant 3 should do the trick

OpenStudy (fibonaccichick666):

check your angle again, I think the mistake may just be miscopying

OpenStudy (anonymous):

i dont get it

OpenStudy (fibonaccichick666):

ok, so 3/4 is less than one right?

OpenStudy (anonymous):

ooo i meant 5pi/4

OpenStudy (fibonaccichick666):

that would make more sense, let me see if that works :)

OpenStudy (fibonaccichick666):

yep! that is it, see told you I thought it was a type-o :)

OpenStudy (anonymous):

but... i dont know how to put in the answer bc everytime i put it in its wrong

OpenStudy (fibonaccichick666):

what are you typing exactly?

OpenStudy (anonymous):

at first i did exact value. then i answered with pi... still wrong.. unless i am typing in the wrong answer

OpenStudy (freckles):

ok so (x,y) the smallest x was pi/4 the largerst x was 5pi/4 for either one of those the y will have to be pi/2 so the points (x,y) from smallest x to largest x: (pi/4,pi/2) (5pi/4,pi/2)

OpenStudy (freckles):

is that what you have been entering?

OpenStudy (freckles):

if not, take a picture of it and post it here so fibo and i can see it.

OpenStudy (fibonaccichick666):

I think the way freckles said it should work

OpenStudy (anonymous):

wait why? im confused.

OpenStudy (fibonaccichick666):

you need an ordered pair, a set of x and y, not just one or the other

OpenStudy (anonymous):

ok ok i understand